# Multi-modal atmospheric characterization of Pictoris b

### v1

@@ Abstract
Context. The characterization of giant exoplanets such as Pictoris b (hereafter Pic b) is now routinely performed with multiple spectrographs and imagers exploring different spectral bandwidths and resolutions, allowing for atmospheric retrieval of spectra with or without the conservation of the planet spectral continuum. The accounting of data multi-modality in the analysis can provide a more comprehensive determination of the planets physical and chemical properties and inform on their formation history.
@@
Aims. We present the first VLTI observations at R 4 000 of Pic b, the first obtained on an exoplanet with GRAVITY at such resolution. We upgrade the forward modelling code ForMoSA to account for the data multi-modality, including low- medium- and high-resolution spectroscopy both based on direct model-data comparison and the analysis of cross-correlation signals. We use ForMoSA to refine the constraints on the atmospheric properties of the exoplanet and evaluate the sensitivity on the retrieved values to the input dataset.
@@
Methods. We obtain four high signal-to-noise (S/N20) spectra of Pic b at K-band with GRAVITY at R 4 000 conserving both the pseudo-continuum and the pattern of molecular absorptions. We use ForMoSA with four grids of self-consistent forward models (Exo-REM, ATMO, BT-Settl, and Sonora) exploring different , log(g), metallicity, C/O, and 12CO/13CO ratio. We then combine the GRAVITY spectra with published 1-5 µm photometry (NaCo, VisAO, NICI, SPHERE), low- to medium- resolution ( broad band, 0.9-7 µm) spectra, and echelle spectra covering narrower bandwidths ( 100 000, 2.1-5.2 µm).
@@
Results. Sonora and Exo-REM are statistically preferred among all four models, regardless of the dataset used. Exo-REM predicts K and log(g) dex when using only the GRAVITY epochs and K and log(g) dex when incorporating all available datasets. The inclusion of archival data significantly affects all retrieved posteriors. Using all datasets, C/O mostly remains solar (C/O = ) while [M/H] reaches super-solar values (0.500.01). We report the first tentative constrain on the isotopic ratio log(12CO/13CO) = 1.12 in Pic b’s atmosphere; however, this detection remains inconclusive due to telluric residuals affecting both the GRAVITY and SINFONI data. Additionally, we estimate the bolometric luminosity as log(L/L⊙) = -4.01 dex. Using a system age of 233 Myr, this bolometric luminosity and the constraints on the dynamical mass of Pic b we constrain the maximum of heavy-element content of the planet to be on the order of 5% (20-80 ).
@@
Conclusions. The joint access to the pseudo-continuum and molecular lines at K-band as provided by GRAVITY have a significant impact on the retrieved metallicity, possibly owing to the collision-induced absorption driving the continuum shape of K-band. Echelle spectra do not dominate the final fit with respect to lower resolution data covering a broader portion of the spectral energy distribution and the latter keeps encapsulating more robust information on . Future multi-modal framework should include a weighting scheme accounting for the bandwidth and central wavelength of the observations.
@@
Direct imaging has presently unearthed a few dozen of planetary-mass companions111https://exoplanet.eu/ across three decades of semi-major axis, orbiting brown-dwarfs up to massive stars hosts, binaries, constituting hierarchical systems, or nested within belts of debris (e.g., Chauvin et al., 2004; Marois et al., 2010; Kuzuhara et al., 2011; Squicciarini et al., 2022; Chomez et al., 2023). This diverse population is likely the result of multiple formation pathways and a rich set of dynamical evolution mechanisms.
@@
The chemical composition of planet atmospheres has long been regarded as an additional way to constrain the formation and the dynamical evolution processes, starting from our Solar System (e.g., Gautier & Owen, 1989; Owen, 1992; Lecluse et al., 1996; Owen & Encrenaz, 2003). Planets formed in disks should display a wide range of bulk enrichments (Boley et al., 2011; Mordasini et al., 2014) and it remains to be determined whether these are detected in the object’s spectra. The temperature of accreted solids (ice, dust and pebbles) and gas in the disk at the planet formation radius is also expected to leave in addition an imprint on atomic and isotopic ratio (Öberg & Bergin, 2016; Morley et al., 2019; Schneider & Bitsch, 2021a, b; Mollière et al., 2022).
@@
The direct imaging technique allows to capture high-quality spectra of giant exoplanets in a few hours of telescope time. Contrary to transit spectroscopy focusing on close-in and strongly irradiated objects, spectral features are formed over a broad range of pressure (several bars) and better account for the bulk composition of the atmosphere. Low-resolution (20 R) integral field spectrographs fed by extreme adaptive-optics systems such as VLT/SPHERE (Beuzit et al., 2019), Gemini/GPI (Macintosh et al., 2014), Subaru/CHARIS (Groff et al., 2016), or LMIRCam/ALES (Skemer et al., 2015, 2018) are collecting near-infrared (1-5 µm) spectra of known imaged exoplanets below 2” around bright stars, revealing water features of M- and L- type companions (e.g., Greenbaum et al., 2018; Currie et al., 2018; Bonnefoy et al., 2016; Chauvin et al., 2017; Doelman et al., 2022) and methane features on a handful of cooler and least massive T-type exoplanets (Rajan et al., 2017; Mesa et al., 2023; De Rosa et al., 2023). Such spectra offer to determine the bulk physical properties of the objects such as the effective temperature , radius R, and surface gravity log(g), and have confirmed that young mid-L to early-T planets have redder under-luminous spectra than more massive and mature brown dwarfs. These properties can be explained by vigorous vertical mixing and reduced gravitational settling that change the relative abundance of methane and carbon monoxide (Hubeny & Burrows, 2007; Moses et al., 2016) and enhance the cloud opacity. Such clouds were recently detected for the first time in the atmosphere of the low-mass late-L companion VHS1256 b using the JWST (Miles et al., 2023).
@@
Ground and space-based medium-resolution (R 2 000-5 000) integral field spectrographs (VLT/SINFONI and ERIS, Keck/OSIRIS, Gemini/NIFS, JWST/NIRSpec and MIRI) can also provide high-fidelity infrared spectra (1-28 µm) of exoplanets (Konopacky et al., 2013; Petrus et al., 2021; Hoch et al., 2023; Miles et al., 2023; Palma-Bifani et al., 2023; Kiefer et al., 2024). These instruments give access to the planet pseudo-continuum emission while de-blending partly the forests of ro-vibrational absorptions from key molecules (CO, H2O, CH4). They have been used to infer the C/O ratio (Hoch et al., 2023, and ref. therein) and even estimate isotopic ratio (Zhang et al., 2021; Gandhi et al., 2023). Using interferometric techniques, GRAVITY and MATISSE on the VLTI have also demonstrated their capability to extract high S/N K-, L-, and M- band lower-resolution (Rλ = 150-700) spectra (2-5 µm) of exoplanets containing both the continuum and lines from which C/O and metallicity (GRAVITY Collaboration et al., 2019, 2020; Blunt et al., 2023; Nasedkin et al., 2024) can be inferred.
@@
Finally, more recently, single-mode fiber fed echelle spectrographs coupled to adaptive-optics systems (HiRISE at VLT Costes et al., 2024; Denis et al., 2025, KPIC at Keck Delorme et al., 2021, REACH at Subaru Kotani et al., 2020) have started providing continuum-subtracted 1-5 µm spectra of exoplanets at Rλ 100 000 over narrow spectral range. Cross-correlation techniques can then be applied for a joint detection and characterization of the objects, even deeply buried into the stellar halo. The inversion of these data start to provide radial and rotational velocities measurements, as well as metallicity, C/O and isotopic ratio determinations for close-in planets (Snellen et al., 2014; Landman et al., 2024; Parker et al., 2024; Morris et al., 2024; Vigan et al., 2024; Xuan et al., 2024; Zhang et al., 2024).
@@
For more than 40 years now since the discovery of Smith & Terrile (1984), the young ( Myr, Mamajek & Bell, 2014) and nearby (19.8 pc, Gaia Collaboration et al., 2023) Pic system has been a unique laboratory for studying planetary formation, planet-disk interactions as well as the physics of young giant planets. The system is made up of a central star, a dust disk and two known giant planets, b and c (see Table 1 for a more complete description). Pic b was the first planet discovered in the system (Lagrange et al., 2009) and has since been thoroughly examined in numerous independent follow-up studies (e.g. Landman et al., 2024; Worthen et al., 2024; Kiefer et al., 2024. and Houllé et al. 2024, subm., see Table 12). The planet has a semi-major axis of 9.930.03 au and dynamical mass of M inferred from a monitoring of its orbital motion started following the initial discovery (2003 epoch published in 2009), high precision astrometry obtained with GRAVITY, radial velocity measurements of the star and more recently the joint analysis of GAIA and HIPPARCOS measurements (Lagrange et al., 2020; Lacour et al., 2021; Brandt et al., 2021).
@@
Following the firsts photometric measurements of Pic b, Bonnefoy et al. (2013) used 1-5 µm photometry to determined the planet to be a young L-type object with an estimated luminosity of log(L/L⊙)= -3.870.08 dex. Their study and the following ones (Males et al., 2014; Chilcote et al., 2017) used these data compared with evolutionary models to extract the first physical parameter constraints. Assuming an age of Myr, evolutionary models predicts between 1600 K and 1800 K, log(g) 4.0 dex and a R 1.5 .
@@
Forward models predictions on the metallic content of Pic b’s atmosphere show a lot of variation in recent works. GRAVITY Collaboration et al. (2020), using GRAVITY R 500 data, found both super-solar ([M/H] ) and sub-solar ([M/H] ) metallicity when adding or not GPI data to their retrieval. Landman et al. (2024) (using K-band CRIRES+ data) obtained a range of metallicities between and depending on the model they used; a result shared by Worthen et al. (2024) (using MIRI data), also determining a large dispersion of values depending on the considered model ([M/H] between and ). Kiefer et al. (2024) (using SINFONI data) found similar sub-solar values of using Exo-REM. The presence of continuum in the data appears to be impactful on the retrieved [M/H] value. The treatment of clouds in models is also equally important for this object as they can lock significant amount of metallic species (Fe, , Si and O). The main molecules predicted to condensate at chemical equilibrium are enstatite Mg2Si2O6, forsterite Mg2SiO4, and corundum Al2O3.
@@
C/O measurements of Pic b have also changed significantly in recent studies. The sub-solar value of C/O obtained by GRAVITY Collaboration et al. (2020) suggested a formation through core-accretion, with a significant alteration of the envelop composition through the in-fall of planetesimals. Sub-solar values were also found by Landman et al. (2024) and Worthen et al. (2024). However, Kiefer et al. (2024) obtained a solar value of C/O using evolutionary-model informed priors but without continuum, a value better aligned with population studies (Nissen, 2013). This value could still suggest a formation through core-accretion, but with a more moderate planetesimal accretion followed by an outward migration.
@@
To our knowledge, no studies have been carried out on the isotopic content of the object’s atmosphere. This is a difficult task because these absorptions are very subtle, and the slightest systematic error in the models or issues with the data can hinder detectability.
@@
In this paper, we propose to delve in the atmosphere of Pic b using the full compilation of available data and adding the first VLTI/GRAVITY observations at R 4 000 of the planet (and the first ever obtained on an exoplanet with the instrument at such spectral resolution). The multi-modality of the data in hand also motivated us to upgrade our forward modeling code ForMoSA222https://formosa.readthedocs.io/en/latest/ (Forward Modeling tool for Spectral Analysis). Our different datasets are presented in Sect. 2. We present the multi-modal version of our Bayesian inference tool ForMoSA in Sect. 3. We analyse with our unified framework each individual epoch of GRAVITY spectra to account for the orbital motion of the planet in Sect. 4 while performing an inversion of the whole set of existing data in Sect. 5 and 6. We discuss our findings and conclusions in Sect. 7 and 8. Additional material is provided in Appendices (Sect. References).
@@
2 Observation and data reduction
This section presents the various spectroscopic and photometric data we use for our analysis, all combined and summarized in Table 2. We will briefly describe the GRAVITY instrument, the observing conditions and the additional post-processing we applied in the following paragraphs:
@@ 2.1 New GRAVITY observations at high-spectral resolution
GRAVITY (GRAVITY Collaboration et al., 2017) is an interferometric spectro-imager that combines light coming from the four 8m Unit Telescopes (UTs) of the VLT. It operates at K-band between 1.97-2.4 µm. The four new epochs of GRAVITY observations of Pic b where obtained in the span of 3 years (2019/11/09, 2019/11/11, 2021/08/27 and 2022/01/24; respective ESO program IDs : 1104.C-0651(A), 1104.C-0651(D) and 1104.C-0651(B)) using the high spectral resolution mode of the instrument, resulting in an effective spectral resolving power of R 4 000. They represent the first observations of an exoplanet at this resolution with the instrument.

### v2

# Multimodal atmospheric characterization of Pictoris b

@@ Abstract
Context. Characterizations of giant exoplanets such as Pictoris b (hereafter Pic b) are now routinely performed with multiple spectrographs and imagers exploring different spectral bandwidths and resolutions, allowing for atmospheric retrieval of spectra with or without the conservation of the planet spectral continuum. The accounting of data multimodality in the analysis could provide a more comprehensive determination of the planets physical and chemical properties and inform on their formation history.
@@
Aims. We present the first VLTI observations at R 4,000 of Pic b obtained for an exoplanet with GRAVITY at such a high resolution. We upgraded the forward modelling code ForMoSA to account for the data multimodality, including low-, medium-, and high-resolution spectroscopy based on both a direct model-data comparison and an analysis of cross-correlation signals. We used the ForMoSA code to refine the constraints on the atmospheric properties of the exoplanet and evaluated the sensitivity of the retrieved values to the input dataset.
@@
Methods. We obtained four high-signal-to-noise (S/N 20) spectra of Pic b in the K band with GRAVITY at R 4,000 conserving both the pseudo-continuum and the pattern of molecular absorptions. We used ForMoSA with four grids of self-consistent forward models (Exo-REM, ATMO, BT-Settl, and Sonora) to explore different , log(g), metallicity, C/O, and 12CO/13CO ratio values. We then combined the GRAVITY spectra with published 1–5 µm photometry (NaCo, VisAO, NICI, and SPHERE), low-to-medium-resolution ( broadband, 0.9–7 µm) spectra, and echelle spectra covering narrower bandwidths ( 100,000, 2.1–5.2 µm).
@@
Results. Sonora and Exo-REM are statistically preferred among all four models, regardless of the dataset used. Exo-REM predicts K and log(g) dex when using only the GRAVITY epochs, whereas we have K log(g) dex when incorporating all available datasets. The inclusion of archival data significantly affects all retrieved posteriors. When using all datasets, C/O mostly remains solar (), while [M/H] reaches super-solar values (0.50 0.01). We report the first tentative constraint on the isotopic ratio log(12CO/13CO) = 1.12 in Pic b’s atmosphere; however, we note that this detection remains inconclusive due to telluric residuals affecting both the GRAVITY and SINFONI data. Additionally, we estimated the bolometric luminosity as log(L/L⊙) = -4.01 dex. Using a system age of 23 3 Myr, along with this bolometric luminosity and the constraints on the dynamical mass of Pic b, we were able to constrain the maximum of heavy element content of the planet to be on the order of 5% (20–80 ).
@@
Conclusions. The joint access to the pseudo-continuum and molecular lines in the K band provided by GRAVITY have a significant impact on the retrieved metallicity, possibly owing to the collision-induced absorption driving the continuum shape of the K band. The echelle spectra do not dominate the final fit with respect to lower resolution data covering a broader portion of the spectral energy distribution and the latter keeps encapsulating more robust information on . Future multimodal frameworks should include a weighting scheme to account for the bandwidth and central wavelength of the observations.
@@
Direct imaging has unearthed a few dozens of planetary-mass companions111https://exoplanet.eu/ across three decades in terms of the semi-major axis, found to be orbiting brown dwarfs up to massive stars hosts, as well as binaries, constituting hierarchical systems, or nested within belts of debris (e.g., Chauvin et al., 2004; Marois et al., 2010; Kuzuhara et al., 2011; Squicciarini et al., 2022; Chomez et al., 2023). This diverse population is likely the result of multiple formation pathways and a rich set of dynamical evolution mechanisms.
@@
The chemical composition of planet atmospheres has long been regarded as an additional way to constrain the formation and the dynamical evolution processes, starting from our Solar System (e.g., Gautier & Owen, 1989; Owen, 1992; Lecluse et al., 1996; Owen & Encrenaz, 2003). Planets formed in disks are expected to display a wide range of bulk enrichments (Boley et al., 2011; Mordasini et al., 2014) and it remains to be determined whether these are detected in the object’s spectra. The temperature of accreted solids (ice, dust, and pebbles) and gas in the disk at the planet formation radius is also expected to leave in addition an imprint on atomic and isotopic ratio (Öberg & Bergin, 2016; Morley et al., 2019; Schneider & Bitsch, 2021a, b; Mollière et al., 2022).
@@
The direct imaging technique allows to capture high-quality spectra of giant exoplanets in a few hours of telescope time. Contrary to transit spectroscopy focusing on close-in and strongly irradiated objects, spectral features are formed over a broad range of pressure (several bars) and better account for the bulk composition of the atmosphere. Low-resolution (20 R) integral field spectrographs fed by extreme adaptive-optics systems such as VLT/SPHERE (Beuzit et al., 2019), Gemini/GPI (Macintosh et al., 2014), Subaru/CHARIS (Groff et al., 2016), or LMIRCam/ALES (Skemer et al., 2015, 2018) are collecting near-infrared (NIR: 1-5 µm) spectra of known imaged exoplanets below 2” around bright stars, revealing water features of M- and L- type companions (e.g., Greenbaum et al., 2018; Currie et al., 2018; Bonnefoy et al., 2016; Chauvin et al., 2017; Doelman et al., 2022) and methane features on a handful of cooler and least massive T-type exoplanets (Rajan et al., 2017; Mesa et al., 2023; De Rosa et al., 2023). Such spectra offer to determine the bulk physical properties of the objects such as the effective temperature , radius R, and surface gravity log(g). These data have confirmed that young mid-L to early-T planets have redder under-luminous spectra than more massive and mature brown dwarfs. These properties can be explained by vigorous vertical mixing and reduced gravitational settling that change the relative abundance of methane and carbon monoxide (Hubeny & Burrows, 2007; Moses et al., 2016) and enhance the cloud opacity. Such clouds were recently detected for the first time in the atmosphere of the low-mass late-L companion VHS1256 b using the JWST (Miles et al., 2023).
@@
Ground and space-based medium-resolution (R 2,000–5,000) integral field spectrographs (VLT/SINFONI and ERIS, Keck/OSIRIS, Gemini/NIFS, JWST/NIRSpec, and MIRI) can also provide high-fidelity infrared spectra (1–28 µm) of exoplanets (Konopacky et al., 2013; Petrus et al., 2021; Hoch et al., 2023; Miles et al., 2023; Palma-Bifani et al., 2023; Kiefer et al., 2024). These instruments give access to the planet pseudo-continuum emission while de-blending partly the forests of ro-vibrational absorptions from key molecules (CO, H2O, CH4). They have been used to infer the C/O ratio (Hoch et al., 2023, and ref. therein) and even to estimate isotopic ratio (Zhang et al., 2021; Gandhi et al., 2023). Using interferometric techniques, GRAVITY and MATISSE on the VLTI have also demonstrated their capability to extract high S/N K-, L-, and M-band data at lower resolution (Rλ = 150–700) spectra (2–5 µm) for exoplanets containing both the continuum and lines from which C/O and metallicity (GRAVITY Collaboration et al., 2019, 2020; Blunt et al., 2023; Nasedkin et al., 2024) can be inferred.
@@
Finally, more recently, single-mode fiber-fed echelle spectrographs coupled to adaptive optics (AO) systems (HiRISE at VLT Costes et al., 2024; Denis et al., 2025, KPIC at Keck Delorme et al., 2021, REACH at Subaru Kotani et al., 2020) have started providing continuum-subtracted 1–5 µm spectra of exoplanets at Rλ 100,000 over a narrow spectral range. Cross-correlation techniques can then be applied for a joint detection and characterization of the objects, even those deeply buried into the stellar halo. The inversion of these data start to provide radial and rotational velocities measurements, as well as metallicity, C/O and isotopic ratio determinations for close-in planets (Snellen et al., 2014b; Landman et al., 2024; Parker et al., 2024; Morris et al., 2024; Vigan et al., 2024; Xuan et al., 2024; Zhang et al., 2024).
@@
For more than 40 years now since the discovery of Smith & Terrile (1984), the young ( Myr, Mamajek & Bell, 2014) and nearby (19.8 pc, Gaia Collaboration et al., 2023) Pic system has been a unique laboratory for studying planetary formation, planet-disk interactions as well as the physics of young giant planets. The system is made up of a central star, a dust disk and two known giant planets, b and c (see Table 1 for a more complete description). Pic b was the first planet discovered in the system (Lagrange et al., 2009) and it has since been thoroughly examined in numerous independent follow-up studies (e.g., Landman et al., 2024; Worthen et al., 2024; Kiefer et al., 2024; Houllé et al., 2025, see Table 15). The planet has a semi-major axis of 9.93 0.03 au and dynamical mass of M inferred from a monitoring of its orbital motion started following the initial discovery (2003 epoch published in 2009), high-precision astrometry obtained with GRAVITY, radial velocity measurements of the star, and, more recently, the joint analysis of GAIA and HIPPARCOS measurements (Lagrange et al., 2020; Lacour et al., 2021; Brandt et al., 2021).
@@
Following the firsts photometric measurements of Pic b, Bonnefoy et al. (2013) used 1–5 µm photometry to determined the planet to be a young L-type object with an estimated luminosity of log(L/L⊙)= -3.87 0.08 dex. Their study and the following ones (Males et al., 2014; Chilcote et al., 2017) used these data compared with evolutionary models to extract the first physical parameter constraints. Assuming an age of Myr, evolutionary models predicts between 1600 K and 1800 K, log(g) 4.0 dex, and a R 1.5 .
@@
Forward models predictions on the metallic content of Pic b’s atmosphere display a lot of variation in recent works. GRAVITY Collaboration et al. (2020), using GRAVITY R 500 data, found both super-solar ([M/H] ) and sub-solar ([M/H] ) metallicity when adding GPI data to their retrieval, while others have not. Landman et al. (2024) (using K-band CRIRES+ data) obtained a range of metallicities between and depending on the model they used; this result was confirmed by Worthen et al. (2024) (using MIRI data), who also determined a large dispersion of values depending on the considered model ([M/H] between and ). Kiefer et al. (2024) (using SINFONI data) found similar sub-solar values of using Exo-REM. The presence of the continuum in the data appears to be impactful on the retrieved [M/H] value. The treatment of clouds in models is also equally important for this object as they can lock up significant amount of metallic species (Fe, Mg, Si, and O). The main molecules predicted to condensate at chemical equilibrium are enstatite Mg2Si2O6, forsterite Mg2SiO4, and corundum Al2O3.
@@
The C/O measurements of Pic b have also displayed a significant range in recent studies. The sub-solar value of C/O obtained by GRAVITY Collaboration et al. (2020) suggested a formation through core-accretion, with a significant alteration of the envelop composition through the infall of planetesimals. Sub-solar values were also found by Landman et al. (2024) and Worthen et al. (2024). However, Kiefer et al. (2024) obtained a solar value of C/O using evolutionary-model informed priors but without continuum, a value that is better aligned with population studies (Nissen, 2013). This value could still suggest a formation through core-accretion, but with a more moderate planetesimal accretion followed by an outward migration.
@@
To our knowledge, no studies have been carried out on the isotopic content of the object’s atmosphere. This is a difficult task because these absorptions are very subtle and the slightest systematic error in the models or issues with the data can hinder detectability.
@@
In this paper, we propose to delve in the atmosphere of Pic b using the full compilation of available data and adding the first VLTI/GRAVITY observations at R 4,000 of the planet (and the first ever obtained on an exoplanet with the instrument at such spectral resolution). The multimodality of the data in hand also motivated us to upgrade our Forward Modeling tool for Spectral Analysis (ForMoSA222https://formosa.readthedocs.io/en/latest/) code. Our different datasets are presented in Sect. 2. We present the multimodal version of our Bayesian inference tool ForMoSA in Sect. 3. We present the unified framework we used to analyse each individual epoch of GRAVITY spectra to account for the orbital motion of the planet in Sect. 4. We describe our inversion of the whole set of existing data in Sects. 5 and 6. We discuss our findings and conclusions in Sects. 7 and 8. Additional material is provided in the appendices.
@@
and $j$$j$footnotetext: Gaia Collaboration et al. (2023).
2 Observations and data reduction
This section presents the various spectroscopic and photometric data we use for our analysis, all combined and summarized in Table 2. We briefly describe the GRAVITY instrument, the observing conditions, and the additional post-processing we applied in the following.
@@ 2.1 New GRAVITY observations at high-spectral resolution
GRAVITY (GRAVITY Collaboration et al., 2017) is an interferometric spectro-imager that combines light coming from the four 8m Unit Telescopes (UTs) of the VLT. It operates in the K band between 1.97-2.4 µm. The four new epochs of GRAVITY observations of Pic b where obtained in the span of three years (2019/11/09, 2019/11/11, 2021/08/27, and 2022/01/24; respective ESO program IDs : 1104.C-0651(A), 1104.C-0651(D), and 1104.C-0651(B)) using the high spectral resolution mode of the instrument, resulting in an effective spectral resolving power of R 4,000. They represent the first observations of an exoplanet at this resolution with the instrument.